Consider f:R2 to R defined by f(x,y) =(x+y)/(√2) if x=y and f(x,y) =0 otherwise ,show.that fx(0,0) =fy(0,0)=0 and Duf(0,0)=1 ,where ,u=(1/√2,1/√2) Deduce that f is not differentiable at (0,0)


1
Expert's answer
2021-03-31T13:12:03-0400

The directional derivatives are "f_x(0,0)=0'=0" and "f_y(0,0)=0'=0". The derivative in direction "u=(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})" is "D_u(0,0)=\frac{1}{\sqrt{2}}f_x+\frac{1}{\sqrt{2}}f_y=\frac{1}{\sqrt{2}}\cdot\frac{1}{\sqrt{2}}+\frac{1}{\sqrt{2}}\cdot\frac{1}{\sqrt{2}}=1". The derivatives in directions "x", "y" and "u" do not coincide. Therefore, the function is not differentiable at (0,0).